# A FLASH on Blazars

### v2

@@ A FLASH on Blazars
Capturing the Radio Realm of 4FGL Blazars with SKAO Pathfinders
@@
This work investigates the multi-wavelength properties of 165 4FGL blazars from the Fermi-LAT fourth source catalogue, looking for with counterparts in the Australian SKA Pathfinder (ASKAP) First Large Absorption Survey in HI (FLASH) continuum. Using high-resolution data from FLASH and complementary radio datasets, combined with archival Atacama Large Millimeter Array (ALMA) observations, we perform detailed spectral energy distribution (SED) analyses across cm-to-mm wavelengths. Our findings reveal that most blazars exhibit re-triggered peaked spectra, indicative of emission dominated by a single emitting region. Additionally, we identify strong correlations between radio and gamma-ray luminosities, highlighting the significant role of relativistic jets in these active galactic nuclei. The inclusion of spectroscopic redshifts from Sloan Digital Sky Survey (SDSS) and Gaia enables a comprehensive analysis of the evolutionary trends and physical characteristics of the sources. Furthermore, we report a tight Radio-X-ray Correlation for Flat Spectrum Radio Quasars, contrasting with the more scattered behaviour observed in BL-Lacs, reflecting their distinct accretion and jet-driving mechanisms. These results provide critical insights into the physics of blazars and their environments, paving the way for future studies with next-generation facilities like the SKA Observatory (SKAO) for radio observations and Cherenkov Telescope Array for gamma-ray studies.
@@ 1 Introduction
The latest radio interferometers are opening up a brand new window on the main characters of the radio sky: the Active Galactic Nuclei (AGN). Observatories like the Murchison Widefield Array (MWA, Tingay et al. 2013) and the Australian SKA Pathfinder (ASKAP) are able to target faint populations at low radio frequencies, efficiently covering wide areas of the sky. The large multi-wavelength statistic of objects surveyed by these telescopes, finally makes it possible to answer several open questions on the physics and the evolutionary role of the different AGN types.
@@
Figure 1: Map showing the coverage of Fermi-LAT 4FGL, with the light pink regions representing the MOC (Multi-Order Coverage) of the FLASH fields utilised in this work.
@@
Within the radio spectrum, the primary emission process is synchrotron radiation, resulting from electrons moving at relativistic speeds as they spiral around magnetic field lines. This process is characteristic of radio AGN, producing the distinctive relativistic jets. From the radio Spectral Energy Distribution (SED), it is possible to retrieve clues about the history of the AGN. For instance, an inverted spectrum is a hint for self-absorption processes, typical of compact sources. On the other end, a break at high radio frequency suggests the presence of ageing processes, while a flat spectrum could be either a hint of efficient energy pumping, continuously refurnishing the jets with energetic electrons, or of the superposition, along the line of sight, of several components. Further, changes in the trend of the SED might be a proxy for reactivation episodes of the AGN. Thus, broad radio spectral coverage is key for reaching a comprehensive view of the radio emission of the AGN.
@@
Among the several classes of AGN, blazars appear rather interesting from a radio point of view. According to the Unified Model by Urry & Padovani (1995), blazars are AGN whose jets are oriented almost ($<$ 15-20 deg) along the line of sight. They are characterised by Doppler boosting of their flux density, apparent superluminal motion, large and rapid variability, and strong, non-thermal emission over the entire electromagnetic spectrum (Padovani 2016).
@@
Figure 2: The illustrations show three sources categorized as ”IN” (Left), ”OUT” (Center), and ”UNCERTAIN” (Right), respectively. These images are based on FLASH maps. The large white circle is aligned with the Fermi-LAT location, sharing the same radius as the Fermi-LAT’s resolution (0.1 degrees). The pink rectangle is positioned at the FLASH continuum island’s location, with dimensions matching those of the island as determined by Selavy (Whiting & Humphreys 2012). The dark green ellipse indicates the size and position of the corresponding RACS-LOW island, and the small white circle marks the GLEAM-X matched source’s location, with a size equivalent to the (median) GLEAM-X resolution (in arcseconds). The pink ”X” and light green diamond denote the best-corresponding positions of the AT20G source to the 4FGL position and the FLASH continuum position, respectively.
@@
Blazars constitute an ideal background target for detecting HI absorption in both the interstellar medium (ISM) and intergalactic medium (IGM), as their intense radio emission enables the detection of absorption at high redshifts, improving our understanding of the distribution and properties of neutral gas in the Universe (e.g., Maina et al. 2022; Srianand et al. 2022). Thus, projects like the First Large Absorption Survey in HI (FLASH, Yoon et al. 2024), made with the ASKAP, leverage the characteristics of blazars to study HI absorption, neutral gas evolution, and jet-environment interactions. The ability of blazars to illuminate cosmologically distant environments makes them indispensable tools for investigating the distribution of neutral gas and its interaction with relativistic jets. Therefore, ahead of the upcoming data from FLASH and the completion of this survey, a first step towards the full accomplishment of using blazars for HI absorption is, indeed, to understand the peculiarities of the blazar population that FLASH is able to target. Understanding the radio properties of blazars in the FLASH continuum catalogue is key for exploiting their potential as astrophysical probes. In particular, their strong gamma-ray emission makes gamma-ray selection an effective way to identify and classify them (Abdo et al. 2010; Nolan et al. 2012; Acero et al. 2015; Giroletti et al. 2016). The Fermi Large Area Telescope Fourth Source Catalog Data Release 4 (4FGL) by Ballet et al. (2023) offers the perfect playground for identifying blazars. Blazars are, in fact, strong gamma-ray emitters through inverse Compton (IC) of the jet electrons with external photons, typically from the disk, synchrotron self-Compton (SSC) from the same electrons producing the radio emission and hadronic processes. Hadronic processes are generated by protons losing energy through synchrotron emission (if the magnetic field is large enough) or photo-meson reactions. In the latter case neutrino emission from the decaying charged pions is also expected (Tavecchio 2017; Padovani et al. 2022b, a). In the 4FGL catalogue, blazars are divided into two subclasses:
  • BL-Lacertae objects (BL-Lac, BLL), characterised by extremely weak or even absent emission lines in their optical spectra. In these objects, gamma-ray emission is predominantly driven by SSC. Their AGN emission is thought to be mainly driven by the rotation of the supermassive black hole.
@@
  • Flat Spectrum Radio Quasars (FSRQ), which exhibit prominent emission lines in their optical spectra. In the gamma-ray regime, their emission is predominantly due to IC. These sources are considered accretion-driven and extremely powerful across the spectrum, displaying the typical flat spectrum in the radio.
@@
The link between these two classes has been highly debated within the framework of the Unified Model. It has been suggested that FSRQ could evolve into BL-Lac as it loses its power, leaving a rotationally driven structure (Cavaliere & D’Elia 2002). Traditionally, all blazars were thought to be characterized by a flat radio spectrum. However, some studies addressing the issue with multi-wavelength data, e.g. (Massardi et al. 2016), have already highlighted that the situation may have varying nuances. Massardi et al. (2016) shows that about 90 % of the blazars in their sample have a remarkably smooth spectrum, well described by a power-law over the full range, suggesting that what dominates in the range from 5 to 150 GHz is a single emitting region, rather than the superposition of different compact regions self-absorbed up to different frequencies. Giroletti et al. (2016) as well draws a scenario that differs from the traditional view, combining MWA Commissioning Survey (MWACS, Hurley-Walker et al. 2014) observations at 180 MHz and gamma-ray emission. They show that extended steep spectrum emission might still be present in quasars at low frequencies, confirming the possible presence of a jet.
@@
Nevertheless, to further explore the correlation between radio and gamma emission, deep and high-resolution radio panchromatic coverage is needed. The advent of wide-area, low-frequency surveys, such as GaLactic and Extragalactic All-sky Murchison Widefield Array eXtended (GLEAM-X, Hurley-Walker et al. 2022) and the ASKAP surveys, have extended the radio coverage of the southern sky to lower frequencies. In particular, the FLASH continuum survey, opens a new window for detailed studies of synchrotron emission and its connection to gamma-ray sources. These data offer an unexplored perspective to the low energy emission of blazars.
@@
In this work, we investigate the radio emission of 4FGL blazars detected in the FLASH continuum survey (Yoon et al. 2024), reconstructing their spectral energy distributions (SEDs) using multi-frequency radio data from FLASH, Rapid ASKAP Continuum Survey (RACS, McConnell et al. 2020; Hale et al. 2021), GLEAM-X (Hurley-Walker et al. 2022), the Australian Telescope compact array 20 GHz (AT20G, Murphy et al. 2010) survey, and the Atacama Large Millimeter Array CALibrators catalogue (ALMACAL,Oteo et al. 2016). We analyse their spectral trends, explore the presence of re-triggered spectral features, and study correlations with gamma-ray and X-ray emissions. Our findings provide new insights into blazar jet properties, their evolutionary states, and their role as cosmological tracers of neutral gas distribution. These results play a fundamental role in projects like FLASH, which leverage the characteristics of blazars to study HI absorption, neutral gas evolution, and jet-environment interactions. This study also lays the groundwork for future investigations with the SKA Observatory (SKAO) and the Cherenkov Telescope Array (CTA), extending blazar studies into the next-generation radio and gamma-ray era.
@@ 2 Sample selection
Table 1: This table lists the surveys used in this study, showing how many counterparts we found and the spectral window covered. 4FGL is the starting point of this selection, counting 3419 blazars candidates.
@@
The starting point for our sample selection was the 4FGL catalogue, DR4 (Ballet et al. 2023). Our sample consists of the BL-Lac (BLLs), FSRQs and Blazar Candidates of Uncertain type (BCUs) from this catalogue with counterparts in the FLASH continuum maps. In particular, we have used the 50 FLASH maps flagged as good, which were downloaded up to November 21, 2024 (Figure 1). Since the survey is still ongoing, the number of publicly available good fields may have increased since then. In this section, we will outline the primary catalogues used and detail the criteria for selecting the sources for our sample. The surveys used in this works are summarised in Table 1.
@@ 2.1 FLASH
FLASH is a radio survey conducted with the ASKAP. FLASH is designed to detect neutral hydrogen in absorption against bright radio continuum sources, probing the diffuse and dense phases of the intergalactic and interstellar medium over cosmological scales. The survey covers a redshift range of $1.00.4<z<$ 1.0. In this work, we make use of the FLASH continuum observations. FLASH continuum observations cover the frequency range 711.5–999.5 MHz, with a typical sensitivity of $\sim$0.2 mJy/beam and an angular resolution of $\sim$15”.
@@ 2.3 Selection process
We performed a cross-match between the two catalogues, using a radius of 0.1 deg, corresponding to the Full Width at Half Maximum (FWHM) of LAT, as, between the two instruments, it possesses the least effective angular separation since ASKAP beam size is of $\sim$ 15 arcsec at 855 MHz (FLASH central frequency). Given the huge difference in the resolving power of the surveys, we visually inspected the data to exclude spurious and ambiguous pairings (Figure 2):
  • As FLASH is composed of isolated mosaics (Figure 1), some of the sources happened to be on the edge of the fields, where the reliability drops down;
  • In some cases, more than one source was equally probable to be the counterpart.
In the second case, we did not directly exclude the sources. Instead, we took advantage of other existing radio surveys, namely AT20G (Murphy et al. 2010; Massardi et al. 2011), RACS (McConnell et al. 2020), and GLEAM-X (Hurley-Walker et al. 2022), to decide, via visual inspection, which FLASH source was the most suitable counterpart. Nevertheless, for some objects it was impossible to determine which was the more likely counterpart, and, thus, we classified them as UNCERTAIN. We got a total of 259 cross-matched sources, 165 of which are classified as IN, 33 as OUT, and 80 as UNCERTAIN.
@@ 2.4 The spectroscopic sub-sample
To find the sources’ spectroscopic redshifts, we cross-matched the FLASH positions of the sources in our sample with the Sloan Digital Sky Survey (SDSS) DR17 survey and the Gaia–unWISE Quasar Catalog (QUAIA), using a radius of 5 arcsec. The SDSS covers $\sim$ 14000 deg2 of the sky with a typical photometric sensitivity of 22.5 magnitudes in five optical bands (u, g, r, i, z) and a spatial resolution of $\sim$ 1.4 arcsec, providing spectra of the objects with a resolution of $R\sim$ 2000 at 3600–10400 $\AA$. The SDSS overlaps FLASH in the equatorial area. QUAIA is an all-sky spectroscopic quasar sample. The catalogue draws on the 6,649,162 quasar candidates identified by the Gaia mission that have redshift estimates from the space observatory’s low-resolution blue photometer/red photometer spectra. As a result, we achieved a total of 77 spectroscopically confirmed objects. The median redshift of these sources is 0.77, consistent with the overall FLASH catalogue (3).
Figure 3: The distribution of spectroscopic redshifts from SDSS for sources marked as ”IN” with a spectroscopic counterpart.
Figure 3 illustrates that BL-Lacs are generally at lower redshifts than FSRQs, bolstering the hypothesis of a parent-child relationship between the two populations. However, a larger spectroscopic sample is required to assess this scenario.
@@ 3.1 Radio properties
$\alpha_{\mathrm{retrig}}$0.5
@@
$\alpha_{\mathrm{thin}}$0.6
@@
To fully reconstruct the radio spectral energy distribution of the sources and, thus, retrieve their properties, we used all the radio data available for the objects in our sample.
@@
A key aspect to consider when analyzing multi-frequency data is the significant disparity in angular resolution across different surveys. At lower frequencies (e.g., GLEAM-X at $\sim$72-231 MHz), the resolution is $\sim$45”, while at 20 GHz (AT20G) it improves to a few arcseconds. This resolution mismatch could introduce uncertainties when associating extended radio structures with gamma-ray sources. However, since most blazars are expected to be compact at GHz frequencies, this effect is likely minimal.
@@
We cross-matched the FLASH coordinates from our catalogue with GLEAM-X, RACS, AT20G, and ALMACAL (Oteo et al. 2016), enabling us to achieve comprehensive photometric coverage across radio frequencies, ranging from 72 MHz (GLEAM-X) to approximately 30 GHz (ALMACAL). This section will outline the development of the radio multi-wavelength catalogue, detail the SED-fitting process, and present the results derived from this procedure.
@@
RACS covers the whole sky at $-90<\delta<49$ at $\sim$887.5/943.5 (RACS-low,McConnell et al. 2020; Hale et al. 2021), 1367.5 (RACS-mid, Duchesne et al. 2023) and 1655.5 MHz (RACS-high,Duchesne et al. 2025), reaching a sensitivity of $\sim$0.25 (RACS-low) and 0.2 mJy/beam (RACS-mid, high) with a resolution of $\sim$ 15” (RACS-low), 10” (RACS-mid) and 8” (RACS-high). Thus, RACS-low observations enable us to explore the sub-GHz emission as well. We cross-matched the catalogue using a 15” radius and, as expected, we found that all of our sources have a counterpart in RACS.
@@
GLEAM-X covers an area of 1447 deg2 around the South Galactic Pole region covering $-32.7<\delta<-20.7$ with a median sensitivity 6.35 mJy/beam and a median resolution of 45” (Hurley-Walker et al. 2022). The main advantage of the GLEAM-X survey is that it is composed of twenty frequency bands in the range 72–231 MHz. This interval samples the MHz side of the radio SED, where self-absorption processes occur.The SED fitting procedure, described in the following sub-section, included all the sources with at least 5 photometric detections in the radio bands.
3.2 Radio SED-fitting

### v3

@@ A FLASH on blazars
Capturing the radio realm of 4FGL blazars with SKA Observatory pathfinders
@@
This work investigates the multi-wavelength properties of 165 sources from the Fermi Large Area Telescope Fourth Source Catalog Data Release 4 (4FGL), with the aim of identifying counterparts in the Australian SKA Pathfinder (ASKAP) First Large Absorption Survey in the HI (FLASH) continuum. Using high-resolution data from FLASH, complementary radio datasets, and archival ALMA observations, we performed detailed spectral energy distribution (SED) analyses across centimetre to millimetre wavelengths. Our findings reveal that most blazars exhibit re-triggered peaked spectra, which are indicative of emission dominated by a single emitting region. Additionally, we identify strong correlations between radio and gamma-ray luminosities, highlighting the significant role of relativistic jets in these active galactic nuclei. The inclusion of spectroscopic redshifts from the Sloan Digital Sky Survey and Gaia enabled a comprehensive analysis of the evolutionary trends and physical characteristics of the sources to be performed. Furthermore, we report a tight radio–X-ray correlation for flat spectrum radio quasars, thus contrasting with the more scattered behaviour observed in BL Lacertae objects and reflecting the distinct accretion and jet-driving mechanisms of the two populations. These results provide critical insights into the physics of blazars and their environments, paving the way for future studies with next-generation facilities such as the SKA Observatory for radio observations and the Cherenkov Telescope Array for gamma-ray studies.
@@ 1 Introduction
The latest radio interferometers are opening up a brand new perspective on the main characters of the radio sky: active galactic nuclei (AGN). Observatories such as the Murchison Widefield Array (MWA; Tingay et al. 2013) and the Australian SKA Pathfinder (ASKAP) are able to target faint populations at low radio frequencies, efficiently covering wide areas of the sky. The large multi-wavelength statistics of the objects surveyed by these telescopes finally makes it possible to answer several open questions on the physics and the evolutionary role of the different AGNs types.
@@
Figure 1: Map showing the coverage of Fermi-LAT 4FGL. The light pink regions represent the Multi-Order Coverage of the FLASH fields utilised in this work.
@@
Within the radio spectrum, the primary emission process is synchrotron radiation, which results from electrons moving at relativistic speeds as they spiral around magnetic field lines. This process is characteristic of radio AGNs, and it produces their distinctive relativistic jets. Through the analysis of the radio spectral energy distribution (SED) of AGNs, it is possible to retrieve clues about their history. For instance, an inverted spectrum is an indication of self-absorption processes, which are typical of compact sources. On the other end, a break at high radio frequency suggests the presence of ageing processes, while a flat spectrum could either be an indication of efficient energy pumping continuously refurnishing the jets with energetic electrons or of the superposition along the line of sight of several components. Further, changes in the trend of the SED might be a proxy for reactivation episodes of the AGN. Thus, broad radio spectral coverage is key to reaching a comprehensive view of the radio emission of AGNs.
@@
Among the several classes of AGNs, blazars appear rather interesting from a radio point of view. According to the Unified Model by Urry & Padovani (1995), blazars are AGNs whose jets are oriented almost (<< 15-20 deg) along the line of sight. They are characterised by Doppler boosting of their flux density, apparent superluminal motion, large and rapid variability, and strong non-thermal emission over the entire electromagnetic spectrum (Padovani 2016).
@@
Figure 2: Illustrations showing three sources categorised as ‘IN’ (Left),‘OUT’ (Center), and ‘UNCERTAIN’ (Right), respectively. These images are based on FLASH maps. The large white circle is aligned with the Fermi-LAT location, sharing the same radius as the Fermi-LAT’s resolution (0.1 degrees). The pink rectangle is positioned at the FLASH continuum island’s location, with dimensions matching those of the island as determined by Selavy (Whiting & Humphreys 2012). The dark green ellipse indicates the size and position of the corresponding RACS-LOW island, and the small white circle marks the GLEAM-X matched source’s location, with a size equivalent to the (median) GLEAM-X resolution (in arcseconds). The pink ‘X’ and light green diamond denote the best-corresponding positions of the AT20G source to the 4FGL position and the FLASH continuum position, respectively.
@@
Blazars constitute an ideal background target for detecting HI absorption in both the interstellar medium (ISM) and intergalactic medium, as their intense radio emission enables the detection of absorption at high redshifts, which can help improve our understanding of the distribution and properties of neutral gas in the Universe (e.g. Maina et al. 2022; Srianand et al. 2022). Thus, projects such as the First Large Absorption Survey in HI (FLASH; Yoon et al. 2025), made with ASKAP, leverage the characteristics of blazars to study HI absorption, neutral gas evolution, and jet-environment interactions. The ability of blazars to illuminate cosmologically distant environments makes them indispensable tools for investigating the distribution of neutral gas and its interaction with relativistic jets. Therefore, ahead of the upcoming data from FLASH and the completion of this survey, a first step towards the full accomplishment of using blazars for HI absorption would be to understand the peculiarities of the blazar population that FLASH is able to target. Understanding the radio properties of blazars in the FLASH continuum catalogue is key to exploiting their potential as astrophysical probes. In particular, their strong gamma-ray emission makes gamma-ray selection an effective way to identify and classify them (Abdo et al. 2010; Nolan et al. 2012; Acero et al. 2015; Giroletti et al. 2016). The Fermi Large Area Telescope Fourth Source Catalog Data Release 4 (4FGL) by Ballet et al. (2023) offers the perfect playground for identifying blazars. Blazars are strong gamma-ray emitters through inverse Compton scattering of the jet electrons with external photons, typically from the disc. They can also produce gamma-rays via synchrotron self-Compton emission from the same electrons responsible for the radio emission, and through hadronic processes. Hadronic processes are generated by protons losing energy through synchrotron emission (if the magnetic field is large enough) or photo-meson reactions. In the latter case, neutrino emission from the decaying charged pions is also expected (Tavecchio 2017; Padovani et al. 2022b, a).
@@
In the 4FGL catalogue, blazars are divided into two subclasses. One of the classes is BL Lacertae objects (BL Lac), which are characterised by extremely weak or even absent emission lines in their optical spectra. In these objects, gamma-ray emission is predominantly driven by synchrotron self-Comptonisation. Their AGN emission is thought to be mainly driven by the rotation of the supermassive black hole. The second class is flat spectrum radio quasars (FSRQs), which exhibit prominent emission lines in their optical spectra. In the gamma-ray regime, their emission is predominantly due to inverse Comptonisation. These sources are considered to be accretion driven and extremely powerful across the spectrum, displaying the typical flat spectrum in the radio.
@@
The link between these two classes has been highly debated within the framework of the Unified Model. It has been suggested that an FSRQ could evolve into a BL Lac as it loses its power, leaving a rotationally driven structure (Cavaliere & D’Elia 2002). Traditionally, all blazars were thought to be characterised by a flat radio spectrum. However, some studies addressing the issue with multi-wavelength data (e.g. (Massardi et al. 2016)) have already highlighted that the situation may have varying nuances. Massardi et al. (2016) showed that about 90% of the blazars in their sample have a remarkably smooth spectrum well described by a power law over the range 5–150 GHz. This suggests that the emission in this frequency interval is dominated by a single emitting region rather than by the superposition of different compact regions self-absorbed up to different frequencies. Giroletti et al. (2016) drew a scenario that differs from the traditional view by combining MWA Commissioning Survey (Hurley-Walker et al. 2014) observations at 180 MHz and gamma-ray emission. They showed that extended steep spectrum emission might still be present in quasars at low frequencies, thus confirming the possible presence of a jet.
@@
Nevertheless, to further explore the correlation between radio and gamma-ray emission, deep high-resolution radio panchromatic coverage is needed. The advent of wide-area low-frequency surveys, such as the GaLactic and Extragalactic All-sky Murchison Widefield Array eXtended (GLEAM-X, Hurley-Walker et al. 2022) and the ASKAP surveys, have extended the radio coverage of the southern sky to lower frequencies. In particular, the FLASH continuum survey is significant in the endeavour to conduct detailed studies of synchrotron emission and its connection to gamma-ray sources, as the data from this survey offer an unexplored perspective on the low-energy emission of blazars.
@@
In this work, we investigate the radio emission of 4FGL blazars detected in the FLASH continuum survey (Yoon et al. 2025), reconstructing their SEDs using multi-frequency radio data from FLASH, the Rapid ASKAP Continuum Survey (RACS; McConnell et al. 2020; Hale et al. 2021), GLEAM-X (Hurley-Walker et al. 2022), the Australian Telescope compact array 20 GHz (AT20G; Murphy et al. 2010) survey, and the Atacama Large Millimeter Array CALibrators catalogue (ALMACAL;Oteo et al. 2016). We analyse their spectral trends, explore the presence of re-triggered spectral features, and study correlations with gamma-ray and X-ray emissions. Our findings provide new insights into blazar jet properties, their evolutionary states, and their role as cosmological tracers of neutral gas distribution. These results will play a fundamental role in projects such as FLASH that leverage the characteristics of blazars to study HI absorption, neutral gas evolution, and jet-environment interactions. This study also lays the groundwork for future investigations with the SKA Observatory and the Cherenkov Telescope Array, extending blazar studies into the next-generation radio and gamma-ray era.
@@ 2 Sample selection
Table 1: Surveys used in this study.
@@
The starting point for our sample selection was the 4FGL catalogue, DR4 (Ballet et al. 2023). Our sample consists of the BL Lac, FSRQs, and blazar candidates of uncertain type (BCUs) from this catalogue with counterparts in the FLASH continuum maps. In particular, we used the 50 FLASH maps flagged as ‘good’, which were downloaded up to November 21, 2024 (Figure 1). Since the survey is still ongoing, the number of publicly available ‘good; fields may have increased since then. In this section, we outline the primary catalogues used and detail the criteria for selecting the sources for our sample. The surveys used in this work are listed in Table 1.
@@ 2.1 FLASH
FLASH is a radio survey conducted with ASKAP (Yoon et al. 2025). It is designed to detect neutral hydrogen in absorption against bright radio continuum sources, and probes the diffuse and dense phases of the intergalactic and ISM over cosmological scales. The survey covers a redshift range of $0.4<z<1.0$. In this work, we make use of the FLASH continuum observations. FLASH continuum observations cover the frequency range 711.5–999.5 MHz, with a typical sensitivity of $\sim$0.2 mJy/beam and an angular resolution of $\sim$15”.
@@ 2.3 Selection process
We performed a cross-match between the two catalogues using a radius of 0.1 deg, which corresponds to the full width at half maximum of LAT, as between the two instruments, LAT possesses the least effective angular separation since the ASKAP beam size is of $\sim$ 15 arcsec at 855 MHz (FLASH central frequency). Given the large difference in the resolving power of the surveys, we visually inspected the data to exclude spurious and ambiguous pairings (Figure 2).
As FLASH is composed of isolated mosaics (Figure 1), some of the sources happened to be on the edge of the fields, where the reliability drops down. In some cases, more than one FLASHsource was equally probable to be the counterpart of one FERMI source. In the second case, we did not directly exclude the sources. Instead, we took advantage of other existing radio surveys, namely AT20G (Murphy et al. 2010; Massardi et al. 2011b), RACS (McConnell et al. 2020), and GLEAM-X (Hurley-Walker et al. 2022), to decide, via visual inspection, which FLASH source was the most suitable counterpart. Nevertheless, for some objects it was impossible to determine which was the more likely counterpart, and thus we classified them as ‘UNCERTAIN’. We obtained a total of 259 cross-matched sources, and of these, we classified 278 as IN, 33 as OUT, and 80 as UNCERTAIN.
@@ 2.4 The spectroscopic sub-sample
[…]
Figure 3: Distribution of spectroscopic redshifts from SDSS for sources marked as ’IN’ with a spectroscopic counterpart.
Figure 3 illustrates that BL Lacs are generally at lower redshifts than FSRQs, bolstering the hypothesis of a parent-child relationship between the two populations. However, a larger spectroscopic sample is required to assess this scenario.
@@ 3.1 Radio properties
$\alpha_{\mathrm{retrig}}$ -0.5
@@
$\alpha_{\mathrm{thin}}$ -0.6
@@
To fully reconstruct the radio SED of the sources and thus retrieve their properties, we used all the radio data available for the objects in our sample. A key aspect to consider when analysing multi-frequency data is the significant disparity in angular resolution across different surveys. At lower frequencies (e.g. GLEAM-X at $\sim$72-231 MHz), the resolution is $\sim$45”, while at 20 GHz (AT20G), it improves to a few arcseconds. This resolution mismatch could introduce uncertainties when associating extended radio structures with gamma-ray sources. However, since most blazars are expected to be compact at GHz frequencies, this effect is likely minimal.
@@
We cross-matched the FLASH coordinates from our catalogue with GLEAM-X, RACS, AT20G, and ALMACAL (Oteo et al. 2016), enabling us to achieve comprehensive photometric coverage across radio frequencies ranging from 72 MHz (GLEAM-X) to approximately 30 GHz (ALMACAL). In this section, we outline the development of the radio multi-wavelength catalogue, detail the SED-fitting process, and present the results derived from this procedure.
@@
The survey RACS covers the whole sky at $90<\delta<49$ at $\sim$887.5/943.5 (RACS-low,McConnell et al. 2020; Hale et al. 2021); 1367.5 (RACS-mid, Duchesne et al. 2023); and 1655.5 MHz (RACS-high,Duchesne et al. 2025), reaching a sensitivity of $\sim$0.25 (RACS-low) and 0.2 mJy/beam (RACS-mid, high) with a resolution of $\sim$ 15” (RACS-low), 10” (RACS-mid), and 8” (RACS-high). Thus, RACS-low observations allowed us to explore the sub-gigahertz emission as well. We cross-matched the catalogue using a 15” radius, and as expected, we found that all of our sources have a counterpart in RACS.
@@
The GLEAM-X survey covers an area of 1447 deg2 around the South Galactic Pole region covering $32.7<\delta<-20.7$ with a median sensitivity of 6.35 mJy/beam and a median resolution of 45” (Hurley-Walker et al. 2022). The main advantage of GLEAM-X is that it is composed of twenty frequency bands in the range 72–231 MHz. This interval samples the megahertz side of the radio SED, where self-absorption processes occur.
Regarding AT20G, it is a blind 20 GHz survey of the entire southern sky conducted with the Australian Telescope Compact Array (ATCA). It reaches a sensitivity of 40 mJy/beam (Murphy et al. 2010) and a resolution of $\sim$ 0.15”. Further, most sources, $\delta<-15.0$, also have near-simultaneous measurements at 5 and 8 GHz. The primary advantage of AT20G for blazar SED studies is its high-frequency coverage and sensitivity over a large sky area (Massardi et al. 2011b).
The ALMACAL survey is constructed from reprocessed ALMA calibration scans suitable for scientific analysis (Oteo et al. 2016). Thus, the fields are randomly distributed across the ALMA-visible sky, cumulatively covering $\sim$ 0.3 deg2. Its frequency coverage spans from 95 to 350 GHz, making this catalogue extremely useful in the SED modelling of bright AGNs.
Variability and temporal aspects of the data
Blazars are intrinsically variable sources across the entire electromagnetic spectrum, including the radio band. Since our spectral analysis is based on multi-frequency archival data, we acknowledge that the non-simultaneity of observations may introduce uncertainties in the derived spectral shapes. However, most of our results—and the sample selection itself—are based on the very low frequency bands, where radio variability is estimated to remain below 5% (Massardi et al. 2011a). At higher frequencies, variability increases significantly, but the lack of recent and deep wide-area millimetric surveys makes it impossible to obtain fully coeval observations. To account for this, we systematically increased the flux uncertainties for the AT20G and ALMACAL measurements. In the following, we provide a brief overview of the time sampling characteristics of the datasets we used to assess the potential impact of variability.
GLEAM-X (Hurley-Walker et al. 2014) observed the 72–231 MHz range with the Murchison Widefield Array between 2018 and 2020. The band was covered in five subbands, cycled through every $\sim$10 minutes, resulting in quasi-simultaneous coverage of the full range for each source. RACS (McConnell et al. 2020) was conducted in separate campaigns for different frequencies: RACS-low (887.5 MHz) from April 2019 to mid 2020 and RACS-mid (1367.5 MHz) in a later epoch. Although the observations are not simultaneous, many sources have data in both bands within a temporal gap of $\lesssim$1.5 years.
FLASH (Yoon et al. 2025) provides continuum fluxes in the 711.5–999.5 MHz range, observed simultaneously within each ASKAP pointing. For sources where both FLASH and RACS data are available, we used them cautiously, being aware of possible offsets due to variability.
AT20G (Murphy et al. 2010; Massardi et al. 2011b) was carried out between 2004 and 2008. Although the initial blind scan was done at 20 GHz, follow-up observations at 5 and 8 GHz were performed within a short time (weeks) of detection, offering a relatively consistent three-frequency snapshot.
ALMACAL (Oteo et al. 2016) spans from ALMA Cycle 0 (2011) to the present. While individual calibrator scans are short and of a single frequency, repeated observations across multiple cycles provide multi-frequency coverage. In this work, we used only fluxes taken within the same or adjacent ALMA cycles to limit the impact of long-term variability.
In summary, although our dataset is heterogeneous and was not selected to minimise variability, in many cases the time intervals between observations are sufficiently short to reduce the most significant inconsistencies. We discuss the possible implications of variability in the interpretation of our results in Section 4.
3.2 Radio SED fitting